I have to multiply using the vertical method. Can I get some assistance and somebody to explain how the process is done?
(2r - 3s)( 4r^2 + 6rs + 9s^2)
Thanks.

Answer by malakumar_kos@yahoo.com(315)


(2r - 3s)( 4r^2 + 6rs + 9s^2)
For vertical multiplication adopt the principal used for integers
(4r^2+6rs+9s^2)(2r-3s)
(4r^2+6rs+9s^2)(2r) = 8r^3+12r^2s+18rs^2
(4r^2+6rs+9s^2)(-3s) = -12r^2s-18rs^2-27s^3
adding th corresponding like terms , we get
8r^3-27s^3 (12r^2s &-12r^2s is 0 and 18rs^2 & -18rs^2
is 0)
hence (4r^2+6rs+9s^2)(2r-3s) = 8r^3- 27s^3